OpenStudy (jonathan34):

Kay made 32 quarts of apple juice. How many liters of apple juice did Kay make? 1 gallon = 4 quarts 1 gallon = 3.785 liters

OpenStudy (anonymous):

You can do this one using unit substitution. We have quarts in the numerator so we need a convertion wtih quarts in the denominator. We have this but it gives us gallons in the numerator so we need another conversion with gallons in the denomnator to cancel. Wwe have this and it gives us liters in the numerator which is what we want. \[32qrt \times \frac{1 gal}{4 qrt} \times \frac{3.785 L}{11 gal}\]

OpenStudy (anonymous):

Sorry that's a typo. Should be ONE gallon at the end not ELEVEN.